OpenStudy (anonymous):

Please help! Solve: e^(2x+1)=50

OpenStudy (anonymous):

log of both sides: 2x+1 = ln(50) x = (ln(50) - 1)/2

OpenStudy (vijay):

1.456

OpenStudy (anonymous):

ln(50) = 3.91202

OpenStudy (anonymous):

Tonks, I have immediate respect for you on the basis of your username. Long live Harry Potter! (Lord of the Rings is still better though)